Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-057_2000-063

Mission Operations Days: 2000/057 to 2000/063
February 26, 2000 00:00:00 GMT to March 03, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA Off; B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BDNominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67Nominal

This Week’s Completed MODIS Activities:

Saturday, February 26, 2000

2000/057 20:30 Real-time - - Turn off BB

Sunday, February 27, 2000

2000/058 02:04:11 ATC – OA15: SD/SDSM Open Double
2000/058 08:39:42 ATC – OA16: SD/SDSM Screened Double
2000/058 11:50 – 12:40 DATA LOSS – unknown downlink reasons. Non-recoverable.
2000/058 14:44:10 Real-time - - Set Itwk/Vdet to 110/220
2000/058 14:45:26 Real-time - - Tun on BB and set to 315K
2000/058 15:15:13 ATC – OA16: SD/SDSM Screened Double
2000/058 18:03 Real-time - - Turn off BB
2000/058 21:50:44 ATC – OA15: SD/SDSM Open Double

Monday, February 28, 2000

2000/059 04:26:15 ATC – OA16: SD/SDSM Screened Double
2000/059 11:01:46 ATC – OA15: SD/SDSM Open Double
2000/059 12:00:00 ATC – SRCA Full Radiometric
2000/059 12:40:39 ATC – OA16: SD/SDSM Screened Double
2000/059 13:48 Real-time – Reset Itwk/Vdet to nominal values (110/226)
2000/059 13:50 Real-time – Turn on BB
2000/059 13:51 Real-time – Set BB to 290K
2000/059 15:02 Real-time – Set BB duty cycle to 33%
2000/059 19:08 Real-time – SMIR VDET Sweep, Itwk = 190, Vdet = 159, 10 steps of 8
2000/059 19:17 Real-time – Reset SMIR Itwk to 110 and Vdet to 226
2000/059 19:43 Real-time – SMIR VDET Sweep, Itwk = 210, Vdet = 159, 10 steps of 8
2000/059 19:53 Real-time – Reset SMIR Itwk to 110 and Vdet to 226
2000/059 20:42 Real-time – SMIR VDET Sweep, Itwk = 230, Vdet = 159, 10 steps of 8
2000/059 20:51 Real-time – Reset SMIR Itwk to 110 and Vdet to 226
2000/059 22:01 Real-time – SMIR VDET Sweep, Itwk = 250, Vdet = 159, 10 steps of 8
2000/059 22:10 Real-time – Reset SMIR Itwk to 110 and Vdet to 226
2000/059 22:33:56 ATC – OA16: SD/SDSM Screened Double

Tuesday, February 29, 2000

2000/060 03:30:34 ATC – OA16: SD/SDSM Screened Double
2000/060 08:27:12 ATC – OA15: SD/SDSM Open Double
2000/060 12:00:00 ATC – SRCA Full Spatial
2000/060 15:02:43 ATC – OA16: SD/SDSM Screened Double
2000/060 19:59:22 ATC – OA16: SD/SDSM Screened Double

Wednesday, March 1, 2000

2000/061 00:56:00 ATC – OA15: SD/SDSM Open Double
2000/061 05:52:38 ATC – OA16: SD/SDSM Screened Double
2000/061 06:44:56 ATC – SRCA Full Spectral, 30W part I
2000/061 08:39:39 ATC – SRCA Full Spectral, 30W part II
2000/061 10:02:05 ATC – SRCA Full Spectral, 10W part I
2000/061 11:57:24 ATC – SRCA Full Spectral, 10W part II
2000/061 14:07:02 ATC – OA16: SD/SDSM Screened Double
2000/061 20:42:33 ATC – OA15: SD/SDSM Open Double
2000/061 21:00:00 Real-time - - BB Duty Cycle to 100%
2000/061 21:30:00 ATC – BB On to 315K

Thursday, March 2, 2000

2000/062 05:10:56 – 05:16:04 – Sector Rotation of heated BB (DCR Off)
2000/062 05:30:00 ATC – BB Off
2000/062 06:35:47 ATC – OA16: SD/SDSM Screened Double
2000/062 09:53:33 ATC – OA16: SD/SDSM Screened Double
2000/062 16:29:04 ATC – OA15: SD/SDSM Open Double
2000/062 21:25:42 ATC – OA16: SD/SDSM Screened Double
2000/062 22:45 Real-Time - - BB Duty Cycle to 33%
2000/062 23:51:18 Real-Time - - Flight Software Inhibit Group 31 Reset
2000/062 23:58:07 Real-Time - - BB Duty Cycle to 33%
2000/062 23:59:17 Real-Time - - BB Turn On, A Side
2000/062 23:58:07 Real-Time - - BB Temp Set to 290K (DN = 1287)

Friday, March 3, 2000

2000/063 04:01:12 ATC – OA16: SD/SDSM Screened Double
2000/063 10:36:43 ATC – OA15: SD/SDSM Open Double
2000/063 17:12:14 ATC – OA16: SD/SDSM Screened Double
2000/063 23:47:45 ATC – OA16: SD/SDSM Screened Double

This Week's Scheduled MODIS Activities Not Completed:

Thursday, March 2, 2000

2000/062 21:45 ATC – BB On to 290K
Flight Software Inhibit Group 31 was still set from previous day’s MOD_BB_OFF (it was not reset). This Inhibit Group prevents ATC stored commands from executing for the Black Body.

Upcoming MODIS Events:

Friday, March 3, 2000

2000/063 23:47:45 ATC – OA16: SD/SDSM Screened Double

Saturday, March 4, 2000

2000/064 06:23:16 ATC – OA15: SD/SDSM Open Double
2000/064 12:58:46 ATC – OA16: SD/SDSM Screened Double
2000/064 14:48?? Real-time – Set SMIR Itwk / Vdet to 110 / 224
2000/064 14:48?? Real-time – Set Blackbody duty cycle to 100%
2000/064 14:48?? Real-time – Set Blackbody temperature to 315K
2000/064 19:34:17 ATC – OA16: SD/SDSM Screened Double
2000/064 22:51:55 ATC – Formatter to Day Rate (13:56 earlier than Nadir Term. Crossing) This allows for all Bands to be recorded for the following activities.
2000/064 22:51:57 ATC – PC DC Restore OFF
2000/064 22:51:59 ATC – PV DC Restore OFF
2000/064 22:52:01 ATC – Sector Rotation to –3400
2000/064 22:52:03 ATC – OA16: SD/SDSM Screened Double
2000/064 23:09:58 ATC – Sector Rotation to 0
2000/064 23:10:00 ATC – PV DC Restore ON
2000/064 23:10:02 ATC – PC DC Restore ON
2000/064 23:10:04 ATC – Formatter to Night Rate (to make up for earlier Day extension)
2000/064 23:10:04 ATC – Formatter to Day Rate (Back on track for nominal rate modes)
2000/065 23:28?? Real-time – Set Blackbody temperature to 270K

Sunday, March 5, 2000

2000/065 03:48:50 ATC – OA15: SD/SDSM Open Double
2000/065 08:45:19 ATC – OA16: SD/SDSM Screened Double
2000/065 13:41:56 ATC – OA16: SD/SDSM Screened Double
2000/065 18:38:35 ATC – OA15: SD/SDSM Open Double
2000/065 19:00?? Real-time – Set SMIR Itwk / Vdet to 110 / 228
2000/065 19:00?? Real-time – Set Blackbody temperature to 315K

Monday, March 6, 2000

2000/066 00:00?? Real-time – Set Blackbody temperature to 270K
2000/066 00:00?? Real-time – Set Blackbody duty cycle to 33%
2000/066 06:10:43 ATC – OA16: SD/SDSM Screened Double
2000/066 11:07:21 ATC – OA16: SD/SDSM Screened Double
2000/066 12:02:34 ATC – OA19: SRCA Full Radiometric
2000/066 12:46:14 ATC – OA15: SD/SDSM Open Double
2000/066 19:00?? Real-time – Set SMIR Itwk / Vdet to 110 / 226
2000/066 21:00:37 ATC – OA16: SD/SDSM Screened Double

Tuesday, March 7, 2000

2000/067 03:36:08 ATC – OA15: SD/SDSM Open Double
2000/067 08:32:46 ATC – OA16: SD/SDSM Screened Double
2000/067 10:11:38 ATC – OA16: SD/SDSM Screened Double
2000/067 11:00:00 ATC – OA26: Set Blackbody to 285K
2000/067 13:00:00 ATC – OA26: Set Blackbody to 290K
2000/067 before 15:00 Real-time – Set Blackbody duty cycle to 100%
2000/067 15:00:00 ATC – OA26: Set Blackbody to 295K
2000/067 17:00:00 ATC – OA26: Set Blackbody to 300K
2000/067 19:00?? Real-time – Set Blackbody temperature to 270K
2000/067 19:00?? Real-time – Set Blackbody duty cycle to 33%
2000/067 20:04:56 ATC – OA15: SD/SDSM Open Double
2000/067 ?? Real-time – ECAL sweep

Wed, March 15, 2000 - IRU Yaw and Roll Slews
Thus, March 16, 2000 - FSS Yaw and Roll Slews
Wed, March 22, 2000 - Drag Makeup Maneuver
Fri, March 24, 2000 - MODIS Roll Maneuver (-19 deg)
Sat, March 25, 2000 - MODIS Yaw Maneuvers
Sun, March 26, 2000 - MODIS Yaw Maneuvers
Mon, March 27, 2000 - Inclination Maneuver
Tue, March 28, 2000 - Post-inclination Drag Makeup Maneuver

MODIS Anomalies:

We had a red alarm at 2000/058 09:50:24 for MOD_VR_PVLW_RN5V. This has been a noisy tlm point all along, but a red alarm is new and unusual. This point was red for one cycle (65 sec) and then cleared. Things have been nominal ever since. No activities were executing at that time. Had the red limit safing procedure been run, Macro 0 would have executed resulting in MODIS exiting Science Mode. Macro 0 turns off the scan assembly, timing generator, all OBCs, etc.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open. Functional checkout of MODIS is complete and was successful. All mechanisms behaved nominally.

First light occurred at 2000/055 15:27:09

MODIS Telemetry Trends:

Telemetry is nominal.

Non-MODIS Significant Events:

On 2000-062 (22:00-23:00): A telemetry loss occurred just as a MODIS science data playback was beginning. Initial beliefs for the cause was an infringement on the MMS planning products “gray zone” for contact times. It appears that the attempted contact duration was slightly more aggressive than the S/C to TDRSS connection would allow. Before the next contact, the MODIS SSR buffer filled up, since it was pretty full when the previous contact was lost. In the process of doing an unscheduled science playback, to pick up what the previous failed playback missed, the sync-up with Whitesands did not transpire correctly and the playback was not recorded. By that time, loss of signal occurred for the short pass. At the beginning of the next pass, a successful playback was executed. So the times that there will be no science data for MODIS is: 2000-062-22:11:42.570 to 2000-062-22:54:25.644 (just shy of 43 minutes)

The transition to normal operations continues as the FOT has begun running our morning status meeting as Project Leads begin the handover.

Limited Life Item Status:

All limited life items are well within lifetime ranges. The precise statistics for each item have been received from LMMS/Valley Forge and will be incorporated next week.